# Preference-free Alignment Learning with Regularized Relevance Reward

### v1

@@ 4 Experiments
We demonstrate the efficacy of R as a reliable reward for alignment (Section 4.2), further advantages of R with ablation study (4.3), and its generalizability as a preference-free alignment approach (Section 4.4).
4.1 Common Experimental Setup
We mainly conduct experiments based on two alignment datasets, Anthropic-HH (Bai et al., 2022a) and AlpacaFarm (Dubois et al., 2023). Also, we train three instruction-tuned models using R with corresponding SFT datasets. We perform PPO (Ziegler et al., 2020) training with rewards from R utilizing only () pairs, i.e., no preference triplets. As in Section 2, we employ Contriever (Izacard et al., 2021) for scoring and 777huggingface.co/facebook/contriever. More implementation details are in Appendix C. We evaluate the resulting models on three GPT-4-based benchmarks for alignment evaluation: Vicuna-Bench (Chiang et al., 2023), MT-Bench (Zheng et al., 2023), and AlpacaEval (Li et al., 2023), and one classifier-based safety benchmark, ToxiGen (Hartvigsen et al., 2022). In addition, we include SuperNI (Wang et al., 2022) and FactScore (Min et al., 2023) to measure alignment tax. More evaluation details are in Appendix D.
4.2 Can R improve human preference?
We compare R with open-source RMs in improving human preference, e.g., producing more helpful and harmless responses, based on Anthropic-HH dataset (Bai et al., 2022a).
Baseline RMs
We include the open-source RMs studied in Section 2 as baseline RMs for PPO training (Ziegler et al., 2020). OpenAssistant (OASST) is trained with four public feedback datasets, WebGPT (Nakano et al., 2021), OpenAI Summary (Stiennon et al., 2020), Anthropic-HH (Bai et al., 2022a), and SyntheticGPT 888Dahoas/synthetic-instruct-gptj-pairwise, based on DeBERTa-V3-Large (He et al., 2021). SteamSHP-XL is another open-source RM trained with SHP (Ethayarajh et al., 2023a) and Helpfulness subset of the Anthropic-HH. Unlike other feedback datasets, the SHP consists of human-generated responses from an online community, Reddit. It is based on FLAN-T5-XL (Chung et al., 2022). Also, we train RM baseline based on LLaMA-2-7B (Touvron et al., 2023) using Anthropic-HH and SHP, LLaMA-2-7B-RM. We include UltraRM-13B, the most capable open-source RM trained on UltraFeedback (Cui et al., 2023). The UltraFeedback is a 64k AI feedback dataset constructed by GPT-4 (OpenAI, 2023). Also, it is further fine-tuned with the public feedback datasets, Anthropic-HH, SHP, and OpenAI Summary. Based on the same experimental setup, we perform PPO training against the reward score by each RM or R.
R is effective for alignment
In Table 3, we can find that our R significantly improves the model’s preference compared to the SFT baseline on all the preference benchmarks, Vicuna, Alpaca, and MT. Surprisingly, R surpasses all the open-source RMs in the Vicuna, MT-Bench, and ToxiGen even though R is a preference-free approach, i.e., zero feedback dataset 999In Appendix F, R outperforms other preference-based learning approaches, including DPO (Rafailov et al., 2023).. As we use only the helpful subset for the PPO, all the PPO models increase their toxicity compared to the SFT baseline. However, our model is found to be the safest among the PPO models, indicating our improvements on the preference benchmarks are achieved while retaining safety. We presume our lower performance in AlpacaEval compared to UltraRM is because AlpacaEval includes neither the reference answer nor the bi-positional judgments, unlike Vicuna-Bench and MT-Bench. On the other hand, we find that PPO training with R increases reward score by UltraRM as well, as shown in Figure 11 of Appendix E.1.
Model AlpacaEval % Win / % Rel SuperNI ROUGE-L FactScore Score / # Facts SFT (AlpacaFarm) 23.2 / 88.2 34.9 38.4 / 19.9 + PPO w. UltraRM 34.1 / 85.1 28.8 33.6 / 25.8 w. R 37.2 / 86.4 34.9 36.8 / 21.4 w. random 34.8 / 82.6 28.4 36.7 / 22.1 w. learned 38.4 / 85.9 31.2 36.4 / 23.1 w. R () 41.5 / 84.8 22.6 33.3 / 42.4 w. 41.6 / 81.3 13.1 32.4 / 39.6 w. 24.1 / 64.5 7.4 33.9 / 56.6
R increases preference while keeping query relevance
We further investigate the results of AlpacaEval, focusing on the difference between response distributions. In particular, we report Relevant Sentence Ratio (% Rel) of models’ responses to compare their distribution related to each component of R. We measure the % Rel by asking whether each sentence in the generated response is relevant to the given query to GPT-4 (OpenAI, 2023). We request GPT-4 to refer to each sentence first and then judge its query relevance as either ‘Relevant’ or ‘Irrelevant’. That is, it indicates sentence-level query relevance of the responses. The regarding prompt for this evaluation is found in Appendix D.2. We measure % Rel of the models based on the AlpacaEval set. As a result, R results in a more relevant response (85.1%) than all the open-source RMs, including UltraRM (83.1%), trained with AI and human preferences 101010More detailed results are found in Table 8.. We visualize their differences in sentence-level relevance in Figure 3. It exhibits scatter plots of the % Rel according to sentence order. We find all the models tend to show a high relevance at the starting point of each response. However, both PPO models (trained with either UltraRM or R) produce irrelevant sentences as the response gets longer. When comparing the two PPO models, the model trained with UltraRM tends to generate longer off-topic content than R. Also, we find that an ensemble of rewards by R and UltraRM further improves the relevance in Appendix E.2.
4.3 Is the reward design of R effective?
We validate R’s reward design choice by ablating each component one by one from the R. We conduct these experiments based on another alignment dataset, AlpacaFarm (Dubois et al., 2023). We further evaluate models on SuperNI (Wang et al., 2022) and Factscore (Min et al., 2023), containing many closed-ended (CE) queries. We intend to capture alignment tax (Ouyang et al., 2022), the trade-offs for the increased preference from the evaluations.
Backbone Model Alignment Method Feedback (Type / #) AlpacaEval % Win / % Rel MT-Bench 1st / 2nd / Avg ToxiGen () SuperNI ROUGE-L Llama-2-7B Llama-2-7B-Chat PPO Human / 1.4M 71.37 / 86.7 6.13 / 5.81 / 5.97 0.0 7.7 Vicuna-7B-v1.5 SFT - / 0 73.72 / 91.5 6.76 / 5.38 / 6.07 0.3 34.3 + R (Ours) PPO / LoRA - / 0 79.13 / 84.4 7.02 / 5.86 / 6.44 0.1 25.8 Mistral-7B Zephyr-beta DPO AI / 62k 87.80 / 92.2 7.65 / 6.75 / 7.20 0.5 13.5 Mistral-H4-SFT SFT - / 0 75.34 / 91.6 6.88 / 6.11 / 6.49 1.5 22.4 + R (Ours) PPO / LoRA - / 0 77.33 / 90.8 7.20 / 6.06 / 6.63 1.4 15.7 Llama-2-13B Llama-2-13B-Chat PPO Human / 1.4M 78.11 / 95.5 6.96 / 6.24 / 6.60 0.0 7.1 Vicuna-13B-v1.5 SFT - / 0 80.16 / 92.1 6.75 / 6.39 / 6.57 0.2 36.0 + R (Ours) PPO / LoRA - / 0 85.14 / 92.6 7.26 / 6.62 / 6.94 0.1 17.4
R pays less alignment tax
In Table 4, we can find R has an advantage in alignment tax (Ouyang et al., 2022), a phenomenon in which the improved preference accompanies degraded performances on other NLP tasks. For example, R reliably improves the preference (% Win) on the AlpacaEval while keeping high relevance (% Rel), zero-shot instruction-following ability (SuperNI), and factuality (FactScore) simultaneously. However, UltraRM significantly sacrifices the performances in the SuperNI and FactScore compared to the SFT baseline, resulting in lower % Win and % Rel than ours. We compare results according to the choice of relevance model in Appendix G.
Relevance contributes to minimize alignment tax
Moreover, we can observe a clear pattern when excluding the relevance scores from the reward. If we do not use according to query type, i.e., R, it shows even higher preference scores with decreased relevance. However, the achievement accompanies significant performance drops on the SuperNI and FactScore by about 12.3% and 3.5% points, respectively. If the is discarded (), the performance of the relevant sentence ratio and ROUGE-L score of SuperNI drop more drastically. Finally, using only the response length as a reward () improves the preference slightly compared to SFT, as reported in Singhal et al. (2023), but it significantly worsens the % Rel and SuperNI score in return.
Effect of branching by query-type
@@ 4.3 Is the reward design of R effective?
If we use random query type, random , for the R instead of the predicted one by GPT-4, its win rate and relevant sentence ratio are degraded on the AlpacaEval. Also, its instruction-following ability is sacrificed, indicating that reward differentiation based on query types is effective. On the other hand, we show that R’s query type classification by GPT-4 can be effectively replaced with in-house LLM, LLaMA-2-7B (Touvron et al., 2023), i.e., learned 111111The training details of the in-house LLM are in Appendix B.2.. It achieves an even better win rate than the original R with competitive performances for SuperNI and FactScore. Additionally, we compare the models’ diversity and consistency according to the query type in Figure 4. We report the Self-BLEU (Zhu et al., 2018) by sampling 10 responses for each input in the AlpacaEval 121212We perform nucleus sampling (Holtzman et al., 2019).. A high Self-BLEU score indicates a consistent response, whereas a low Self-BLEU score denotes a diverse response (Touvron et al., 2023). As we intended, R improves the consistency in the CE type queries and promotes more diverse responses in the OE type. However, R produces excessively diverse responses regardless of the query type, demonstrating the efficacy of the branching by query type in R.
4.4 Is R generalizable method?
Finally, we compare our R with other state-of-the-art alignment models in various backbones. Note that our method can be adopted upon any instruction-tuned models without preference datasets. To show the generalizability of our method, we conduct PPO training on the three instruction-tuned models, Vicuna-{7, 15}B-v1.5 (Chiang et al., 2023; Zheng et al., 2023) and Mistral-H4-SFT (Tunstall et al., 2023). Specifically, we sample datasets from ShareGPT (Chiang et al., 2023) and UltraChat (Ding et al., 2023), originally used for the SFT training for each instruction-tuned model, for our PPO training. More training details are in Appendix C.
R consistently improves various SFT models
In Table 5, our method consistently improves the aligned behaviors of the three instruction-tuned models while keeping safety and incurring less alignment tax without explicit human or AI feedback datasets. Specifically, our resulting PPO models significantly improve the scores of AlpacaEval (Li et al., 2023) and MT-Bench (Zheng et al., 2023) based on the Vicuna-{7, 15}B-v1.5. They outperform other alignment models within the same backbones, i.e., Llama-2-{7, 13}B-Chat. Our model still increases the preference for the Mistral-H4-SFT backbone, but the increments are behind the Zephyr-beta (Tunstall et al., 2023), which is trained with GPT-4-generated feedback (Cui et al., 2023) 131313Zheng et al. (2023) report that GPT-4 prefers responses from the model trained with GPT-4 outputs, i.e., self-enhancement bias.. We presume the relatively slight improvements on the Mistral-H4-SFT are because the SFT model is less tuned according to the training recipe of Tunstall et al. (2023). On the other hand, we find that relevance (% Rel) is sacrificed according to the increment of preference (% Win) in the 7B models, as in the previous study, but the relevance score is even improved in the 13B model. We speculate the trade-off might be mitigated by scaling the model size. Also, we compare R with other preference-free approaches in Appendix H.
5 Related Work
Preference-based Alignment Learning
Reinforcement Learning from Human Feedback (RLHF) is the most prevalent approach to leverage the human preference (Ziegler et al., 2020; Ouyang et al., 2022; Bai et al., 2022a; Liu et al., 2022; Scheurer et al., 2023). These approaches include the reward modeling stage to develop a proxy RM predicting the human preference over the response, e.g., typically as a scalar format. Then, the trained proxy RM computes reward scores for the sampled responses in the RL stage, i.e., rollout. Recently, replacing human feedback with AI or synthetic feedback has been studied (Bai et al., 2022b; Kim et al., 2023a; Sun et al., 2023b; Cui et al., 2023). Direct Preference Optimization (DPO) reparameterizes the RLHF procedure to the efficient binary cross-entropy objective, eliminating additional RM training and rollouts (Rafailov et al., 2023). Zhao et al. (2023) apply hinge loss between preference pairs in an offline manner. Azar et al. (2023) suggest IPO to mitigate an overfitting problem in the DPO. Ethayarajh et al. (2023b) introduce Kahneman-Tversky Optimization (KTO) requiring only good or bad feedback of responses for given inputs, instead of the pairwise preference. A line of research extends the binary preference to ranking optimization (Yuan et al., 2023; Song et al., 2023a). There are attempts to leverage multiple rewards beyond the overall reward (Touvron et al., 2023; Wu et al., 2023; Rame et al., 2023; Jang et al., 2023). Wu et al. (2023) devise fine-grained rewards consisting of relevance, factuality, and information completeness to enhance long-form question-answering. To this end, human annotators collect the corresponding fine-grained feedback. On the contrary, our study focuses on another alternative to align LLMs leveraging off-the-shelf relevance model (Izacard et al., 2021), while not relying on the explicit preference dataset.
Pitfalls of Reward Modeling and Human Feedback
RLHF faces various challenges, especially in reward modeling and human feedback itself (Casper et al., 2023). Gao et al. (2023) present the overoptimization problem by the imperfect proxy reward models. The issue arises when the policy, tailored to enhance the estimated rewards, fails to improve real-world rewards. Many studies try to mitigate the overoptimization problem (Moskovitz et al., 2023; Coste et al., 2023). Similarly, Shen et al. (2023) study reward inconsistency and leverage a sentence encoder to increase the reward consistency. The work leverages SimCSE (Gao et al., 2021) to augment the preference dataset and normalize reward scores. On the contrary, we directly employ an off-the-shelf retriever, Contriever (Izacard et al., 2021), to compute the rewards. The policy models within the RLHF often find and exploit undesired shortcuts or errors in the RMs to maximize the proxy rewards. The phenomenon is known as reward hacking (Amodei et al., 2016; Pan et al., 2022; Pang et al., 2022; Song et al., 2023b; Eisenstein et al., 2023). On the other hand, much literature points out that human preference can be easily biased by superficial features such as complexity and assertiveness (Xu et al., 2023; Hosking et al., 2023). Singhal et al. (2023) study a high correlation between response length and RLHF training. Motivated by these studies, our work also reveals the poor performance of open-source RMs, trained by human feedback, over trivial relevance problems. Also, the motivation leads us to develop a preference-free approach for aligning models while effectively mitigating reward hacking.

### v3

@@ 4 Experiments
4.1 Reverse Reward Engineering
We compare four combinations of the features as a proxy reward function, (1) LI, (2) , (3) , and (4) RER. To capture the monotonic relationship between proxy and both reward signals, we compute reward scores of both proxy and gold for AlpacaEval Li et al. (2023) set at every 500 PPO steps against the proxy reward. Figure 3 and Table 1 show the results for the reverse reward engineering based on the Anthropic-HH Bai et al. (2022a) and AlpacaFarm Dubois et al. (2023) datasets. We perform standard normalization over the reward scores for the plotting.
We find that the solely promoting lengthy response, i.e., LI as a proxy reward, fails to monotonically increase the gold reward signal, contrary to recent findings in Singhal et al. (2023). More specifically, it improves the gold reward for the initial 1-1.5k steps but subsequently leads to a consistent decrease in the gold reward for the remaining steps even though the proxy rewards are consistently improved, as shown in Figure 3 (a) and (e). Recalling that the PPO steps were set to at most 400 in the Singhal et al. (2023), we presume that the LI is effective only for a few initial steps. Penalizing unnecessary repetitions along with lengthy responses shows a better tendency, yet it eventually results in a decrease in the gold reward (Figure 3 (b) and (f)). However, the magnitude of this decrease is considerably less than that experienced with LI alone. The LI is the only reward that exhibits negative correlations with the gold reward signal in Table 1.
Considering relevance along with the features reliably increases the gold reward, indicating the success of reverse engineering. We find the QR contribute to avoiding the drastic drop in gold reward score and even achieves the success of the reverse engineering, improving the SpearmanR up to 0.86, as shown in Figure 3 (c), (g) and Table 1. On the other hand, RER monotonically increases the gold reward signal for both datasets, as illustrated in Figure 3 (d) and (h). Also, in Table 1, RER achieves the highest Spearman correlations, indicating a strong monotonic relationship with the gold reward signal on both datasets. Although RER shows a relatively slower start compared to other reward options, it exhibits the most consistent upward trend. However, we find the relevance alone, e.g., QR, can not improve the gold reward while showing a similar tendency with the LI. It implies that aiming for sufficiently lengthy and faithful responses while avoiding off-topic or unnecessary repetitions is key to maximizing the gold reward signal. These results have a connection with the previous findings on superficial biases like assertiveness in human preference Xu et al. (2023); Hosking et al. (2023).
Reward branching according to whether the query requires open-ended responses makes a meaningful difference, especially for CE type queries. We compare the final policies optimized and selected by proxy reward functions. Specifically, we measure reward scores using the Gold RM based on the AlpacaEval set, as reported in Figure 4. We find the reward scores are decreased if the model exploits LI regardless of the query type, i.e., no reward branching. We will show this gap makes further differences in the alignment tax problem Askell et al. (2021); Ouyang et al. (2022), which is a phenomenon showing performance drops in other NLP tasks after alignment, in Section 4.3.
@@ 4.1 Reverse Reward Engineering
Additionally, we compare the models’ diversity and consistency according to the query type in Figure 5. We report the Self-BLEU Zhu et al. (2018) by sampling 10 responses for each input in the AlpacaEval 999We perform nucleus sampling Holtzman et al. (2019).. A high Self-BLEU score indicates a consistent response, whereas a low Self-BLEU score denotes a diverse response Touvron et al. (2023). As we intended, RER improves the consistency in the CE type queries and promotes more diverse responses in the OE type, showing a similar tendency to that of Gold RM. However, other models without reward branching produce excessively diverse responses regardless of the query type.
Reward Anthropic-HH AlpacaFarm PearsonR SpearmanR PearsonR SpearmanR w. LI -0.77 -0.66 -0.52 -0.58 w. 0.52 0.05 0.70 0.63 w. 0.96 0.86 0.78 0.86 w. RER 0.92 0.99 0.86 0.97
4.2 Comparison with other open-source RMs
We compare our reward designs with open-source RMs in improving human preference, e.g., producing more helpful responses, based on the Anthropic-HH dataset Bai et al. (2022a).
Model (Anthropic-HH) Vicuna Bench Alpaca Eval MT Bench SFT 50.0 15.9 3.9 + PPO w. SteamSHP-XL 59.4 19.6 3.8 w. OASST 61.3 19.1 3.9 w. UltraRM-13B 73.8 28.0 4.2 w. StarlingRM-34B 79.4 29.1 4.3 w. LI 67.5 16.9 3.4 w. 71.9 20.3 3.9 w. 71.9 23.2 4.0 w. RER 76.9 23.4 4.4
Optimizing only towards lengthy responses, i.e., LI, still improves the alignment scores compared to the SFT baseline, even if it shows drastic drops in Gold RM score as in Section 4.1. In Table 2, the PPO model even performs better SteamSHP-XL and OASST based PPO models in Vicuna-Bench Chiang et al. (2023). This result reminds us of the existence of verbosity bias in LLM-as-a-judge evaluation Zheng et al. (2023). On the other hand, the contributions of RP, QR, and reward branching are shown as a similar pattern with the results of Section 4.1. Notably, the PPO model trained with RER achieves even better performances than the model trained with UltraRM-13B in Vicuna-Bench and MT-Bench.
Model (Anthropic-HH) % Win Rate Avg # Tokens 4-gram Rep. % Rel. Sent SFT 15.9 126 0.05 85.9 + PPO w. SteamSHP-XL 19.6 154 0.10 80.2 w. OAAST 19.1 186 0.10 85.0 w. UltraRM-13B 28.0 213 0.11 83.1 w. StarlingRM-34B 29.1 220 0.08 85.3 w. LI 16.9 351 0.32 76.5 w. 20.3 285 0.17 81.6 w. 23.2 305 0.13 84.8 w. RER 23.4 243 0.09 85.1
PPO model with RER achieves competitive performance while showing a response distribution similar to that of the model with the Gold RM in terms of length, repetitions, and relevance. In Table 3, we report the average number of tokens (Avg #Tokens), 4-gram Repetitions (Rep.), and Relevant Sentence Ratio (% Rel. Sent) of models’ responses based on the AlpacaEval set. We can see the obvious differences in response distribution among RER and other designed reward functions. In particular, it exhibits the importance of reward branching in that it achieves the win rate without unnecessary verbosity, i.e., much less response length and repetition while containing more on-topic content (the higher % Rel. Sent). Qualitative examples among the resulting models are in Appendix F.
Model (AlpacaFarm) AlpacaEval % Win / % Rel SuperNI ROUGE-L SFT 23.2 / 88.2 34.9 + PPO w. UltraRM-13B 34.1 / 85.1 28.8 w. StarlingRM-34B 44.0 / 87.3 17.3 w. LI 24.1 / 64.5 7.4 w. 41.6 / 81.3 13.1 w. 41.5 / 84.8 22.6 w. RER 37.2 / 86.4 34.9 w. random 34.8 / 82.6 28.4
4.3 Investigating Alignment Tax
We further analyze ‘alignment tax’ from the models trained with our devised rewards. The alignment tax is a phenomenon in which the improved preference accompanies degraded performances on other NLP tasks after alignment procedure Askell et al. (2021); Bai et al. (2022a); Ouyang et al. (2022). We conduct these experiments based on another alignment dataset, AlpacaFarm Dubois et al. (2023). Specifically, we evaluate models on SuperNI Wang et al. (2022), the collection of zero-shot NLP tasks, to measure alignment tax.
Reward branching and relevance contribute to minimizing alignment tax.
As shown in Table 4, RER reliably improves the preference (% Win) on the AlpacaEval while keeping high relevance (% Rel), zero-shot instruction-following ability (SuperNI). If we use random query type, random , for the RER instead of the predicted one by GPT-4, its win rate and relevant sentence ratio are degraded on the AlpacaEval. Also, its instruction-following ability is sacrificed, indicating that reward differentiation based on query types is effective in mitigating alignment tax.
There is an observable pattern that clearly exacerbates the alignment tax when excluding each component from RER. If we do not use reference answer relevance according to query type, i.e., , it shows even higher preference scores (% Win) with decreased relevance (% Rel). However, the achievement accompanies significant performance drops on the SuperNI by about 12.3% points. If the QR is discarded (), the performance of the relevant sentence ratio and ROUGE-L score of SuperNI drop more drastically. Finally, using only the response length as a reward (LI) improves the preference slightly compared to SFT, as reported in Singhal et al. (2023), but it significantly worsens the % Rel and SuperNI score in return. On the other hand, while StarlingRM-34B significantly improves the win rate, it notably degrades performance in the SuperNI compared to the SFT baseline, exhibiting a severe alignment tax.
4.4 Generalizability of RER
We investigate how RER is generalizable regarding (1) the number of PPO steps and (2) backbone model size. Specifically, we conduct the experiments based on strong SFT models, Vicuna-{7, 13}B-v1.5 Chiang et al. (2023); Zheng et al. (2023). We sample datasets from ShareGPT Chiang et al. (2023) for our PPO training. More training details are in Appendix C.
We find the monotonic relationship is consistently observed for both backbone sizes even if we extend the number of PPO steps from 5k to 10k, as shown in Figure 6. However, the upward tendency is saturated after the 5k step and often exhibits fluctuation in the reward scores. It appears that such saturation occurs earlier (2k) in the 13B model. Additionally, we observe that RER consistently improves the win rate by approximately 5-6% points for the two SFT backbones in the AlpacaEval, as illustrated in Figure 7. We emphasize that RER could serve as a simple yet robust baseline reward, eliminating the need for a separate reward modeling process that involves pairwise human feedback.
5 Discussion and Future Works
Our study does not cover all perspectives on the role of RMs, such as safety or reasoning. We try to reverse reward engineering over ‘general helpfulness’, one of the perspectives on which many alignment learning studies have focused Li et al. (2023). In other words, we do not argue that our findings are universally applicable. Additionally, the design choice of white-box rewards could be substituted with other features, depending on the target value or domain of interest. Exploring a broader range of interpretable features could enhance our understanding Mu et al. (2024), and investigating aspects of the Gold RM that these features do not capture presents another promising direction for future research.
RM Chat Chat-Hard Safety Reasoning Overall Black-box Rewards UltraRM-13B 97.2 57.0 60.0 73.8 70.6 StarlingRM-34B 97.5 55.7 86.5 90.6 85.1 White-box Rewards LI 77.7 32.0 41.5 57.7 52.2 83.5 32.0 41.1 54.4 51.2 83.5 32.5 41.4 42.0 45.4 RER 80.2 46.3 62.3 57.0 59.5
On the other hand, Table 5 shows the preference accuracy of RMs used in our study. As expected, there is a huge performance gap between our white-box rewards and black-box RMs. However, we observe that the higher performance on the intrinsic evaluation (preference accuracy) does not guarantee a higher alignment performance. For instance, although RER performs worse than the other black-box RMs in preference accuracy, it achieves higher performance on the AlpacaEval and SuperNI than the UltraRM-13B, as we showed in Table 4. We think one possible variable to interpret is that an imperfect reference model is used to compute the KL divergence term Chen et al. (2024a); Ramé et al. (2024). Investigating this misalignment is also an interesting future direction.